# High-dimensional Autoregressive Modeling for Time Series with Hierarchical Structures

### v1

@@ 2.1 Tucker tensor regression and supervised factor modeling
For simplicity, we denote $q_{N+m}=p_{m}$, $s_{N+m}=r_{m}$ and $\boldsymbol{U}_{N+m}=\boldsymbol{V}_{m}$ for $1\leq m\leq M$. Note that the Tucker decomposition at (2.2) is not unique, since $\mbox{$\mathscr{A}$}=\mbox{$\mathscr{G}$}\times_{1}\boldsymbol{U}_{1}\times_{2}\cdots\times_{N+M}\boldsymbol{U}_{N+M}=\left(\mbox{$\mathscr{G}$}\times_{1}\boldsymbol{O}_{1}\times_{2}\cdots\times_{N+M}\boldsymbol{O}_{N+M}\right)\times_{1}\left(\boldsymbol{U}_{1}\boldsymbol{O}_{1}^{-1}\right)\times_{2}\cdots\times_{N+M}\left(\boldsymbol{U}_{N+M}\boldsymbol{O}_{N+M}^{-1}\right)$ for any invertible matrices $\boldsymbol{O}_{i}\in\mathbb{R}^{s_{i}\times s_{i}}$ with $1\leq i\leq N+M$. Without loss of generality, we restrict these factor matrices to be orthonormal, i.e., $\boldsymbol{U}_{j}\in\mathbb{O}^{q_{j}\times s_{j}}$ for each $1\leq j\leq N+M$; see, for example, the higher-order singular value decomposition (HOSVD) in kolda2009tensor. Consequently, the Tucker tensor regression model can be rewritten into
@@
indicating that $\text{col}(\boldsymbol{V})$ contains all useful information of predictors in terms of predicting responses. As a result, define $\boldsymbol{f}_{t}^{\mathrm{response}}=\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t}\times_{n=1}^{N}\boldsymbol{U}_{n}^{\top})=\boldsymbol{U}^{\top}\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})\in\mathbb{R}^{s}$ and $\boldsymbol{f}_{t}^{\mathrm{predictor}}=\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t}\times_{m=1}^{M}\boldsymbol{V}_{m}^{\top})=\boldsymbol{V}^{\top}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})\in\mathbb{R}^{r}$ to be the response and predictor factors, respectively, and then the Tucker tensor regression can be well interpreted from a supervised factor modeling perspective (wang2022high; wang2024high; huang2025supervised).
The above supervised factor modeling interpretation also motivates us to introduce a new framework to construct a tensor regression model, especially for data with complex structures beyond the capability of Tucker decomposition, or for non-orthogonal loadings in the feature extraction process such as the CP decomposition (harshman1970foundations). Specifically, we first extract features or factors from responses and predictors by
@@
respectively. Here, $\mathcal{M}(\cdot):\mathbb{R}^{p_{1}\times\cdots\times p_{M}}\to\mathbb{R}^{r}$ and $\mathcal{N}(\cdot):\mathbb{R}^{q_{1}\times\cdots\times q_{N}}\to\mathbb{R}^{s}$ are feature-extracting operators, with $r$ and $s$ being the number of extracted predictor and response features, respectively. A specific structure is then assumed to the coefficient tensor $\mathscr{A}$ of model (2.1) such that $\boldsymbol{f}_{t}^{\mathrm{response}}=\boldsymbol{\Theta}\boldsymbol{f}_{t}^{\mathrm{predictor}}+\boldsymbol{e}_{t}$, where $\boldsymbol{\Theta}\in\mathbb{R}^{s\times r}$ and $\boldsymbol{e}_{t}\in\mathbb{R}^{s}$ are the corresponding coefficient matrix and compressed error term, respectively. To this end, if operators $\mathcal{M}$ and $\mathcal{N}$ are linear, i.e., $\mathcal{M}(\mbox{$\mathscr{X}$}_{t})=\boldsymbol{\Lambda}_{x}^{\top}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})$ and $\mathcal{N}(\mbox{$\mathscr{Y}$}_{t})=\boldsymbol{\Lambda}_{y}^{\top}\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})$, with non-degenerating loading matrices $\boldsymbol{\Lambda}_{x}\in\mathbb{R}^{P\times r}$ and $\boldsymbol{\Lambda}_{y}\in\mathbb{R}^{Q\times s}$, we then can set the coefficient tensor to
@@ 2.2 Purely hierarchical structure
According to technological advances, the available multi-dimensional or tensor-valued data have become more and more complex, and it is very common to observe hierarchical relationships among some modes. In order to exploit their structural information, we attempt to design a regression model using the proposed framework at (2.4) and (2.5). This subsection starts from the simple case with purely hierarchical responses and predictors.
@@
We first consider the feature extraction on predictors, and assume for simplicity that $M=2$, i.e., $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}}$, and the $p_{1}$ variables of the first mode are nested within each of the $p_{2}$ variables in its second mode. There are two commonly used methods in the literature of factor analysis to extract features from such hierarchical data $\mbox{$\mathscr{X}$}_{t}$: the hierarchical factor model (HFM) in schmid1957development and higher-order factor model (HOFM) in yung1999relationship. The HFM directly extracts features from each level of the hierarchy separately and jointly. In contrast, the HOFM performs feature extraction in a sequential manner along the hierarchy, where features extracted from a lower level are used to summarize those from the upper level. The loading matrices for these two methods are illustrated in Figure 1. Note that the HFM has more than $p_{2}$ features, while the HOFM has $O(rP)$ parameters with $P=p_{1}p_{2}$, making both of them infeasible in high-dimensional settings. Moreover, Tucker decomposition method extracts features simultaneously from all modes, i.e. $\mathcal{M}(\mbox{$\mathscr{X}$}_{t})=\operatorname*{vec}(\boldsymbol{V}_{1}^{\top}\mbox{$\mathscr{X}$}_{t}\boldsymbol{V}_{2})$, and it is not suitable to handle such hierarchical relationships.
@@
Note that the HOFM employs different loadings for each level of the hierarchy. Such flexibility can maintain its performance, while the corresponding number of parameters is also enlarged dramatically for a larger dimension of $M$. This paper alternatively proposes a sequential feature extraction method with shared loadings within each level of the hierarchy. Specifically, we define the operator $\mathcal{M}(\mbox{$\mathscr{X}$}_{t};\{\boldsymbol{G}_{m}\}_{m=1}^{2})=\boldsymbol{G}_{2}^{\top}(\boldsymbol{I}_{p_{2}}\otimes\boldsymbol{G}_{1}^{\top})\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})$, where $\boldsymbol{G}_{1}\in\mathbb{O}^{p_{1}\times r_{1}}$ and $\boldsymbol{G}_{2}\in\mathbb{O}^{p_{2}r_{1}\times r_{2}}$ are the component matrices for summarizing predictors, where $r_{1}$ and $r_{2}$ are the interim dimensions of features extracted from the corresponding mode. To interpret this construction, we liken the hierarchy of $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}}$ to a tree structure with $p_{2}$ branches, each branch containing $p_{1}$ leaves. For each branch, $\boldsymbol{G}_{1}^{\top}$ extracts $r_{1}$ features from its leaves to represent this branch. Overall, we have $r_{1}p_{2}$ features extracted from all leaves, given by $(\boldsymbol{I}_{p_{2}}\otimes\boldsymbol{G}_{1}^{\top})\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})$. Then the subsequent $\boldsymbol{G}_{2}^{\top}\in\mathbb{R}^{r_{2}\times p_{2}r_{1}}$ combines the extracted features from all branches to form $r_{2}$ features for the entire tree; see Figure 1(c) for the illustration of this hierarchical feature extraction procedure. Therefore, the overall loading matrix connecting the $p_{1}p_{2}$ leaves to the final $r_{2}$ features for the tree is $\boldsymbol{G}_{2}^{\top}(\boldsymbol{I}_{p_{2}}\otimes\boldsymbol{G}_{1}^{\top})$, and the column space of this loading can dwell in an arbitrary subspace of $\mathbb{R}^{r_{2}}$.
@@
Consider a special case with $p_{2}=2$, i.e., predictors $\mbox{$\mathscr{X}$}_{t}$ can be divided into two branches, $\boldsymbol{x}_{1:p_{1},1}$ and $\boldsymbol{x}_{1:p_{1},2}$. Suppose that the first branch has common features $\boldsymbol{G}^{*\top}\boldsymbol{x}_{1:p_{1},1}\in\mathbb{R}^{2}$ and branch-specific one $\boldsymbol{G}_{1}^{*\top}\boldsymbol{x}_{1:p_{1},1}\in\mathbb{R}$, while the second one has the common and branch-specific features of $\boldsymbol{G}^{*\top}\boldsymbol{x}_{1:p_{1},2}\in\mathbb{R}^{2}$ and $\boldsymbol{G}_{2}^{*\top}\boldsymbol{x}_{1:p_{1},2}\in\mathbb{R}^{2}$, respectively. Note that the HOFM uses loading matrices $\boldsymbol{G}_{1}^{(1)}\in\mathbb{R}^{p_{1}\times 3}$ and $\boldsymbol{G}_{1}^{(2)}\in\mathbb{R}^{p_{1}\times 4}$ to extract features from the first and second branches, respectively, while the proposed method employs a common loading matrix $\boldsymbol{G}_{1}\in\mathbb{R}^{p_{1}\times 5}$ to extract features from each branch. Shortly speaking, the HOFM concentrates more on branch-specific features, while our method focuses more on common features. In real applications, some weak branch-specific features may be ignored by the introduced method at the first layer of feature extraction, but they may also eventually be averaged out by the HOFM at later layers, i.e., our method does not miss globally important features.
@@
Consider a special case with $r=s$ and $\boldsymbol{\Theta}$ being diagonal, and then the above coefficient tensor $\mathscr{A}$ will admit a form of tensor train decomposition (oseledets2011tensor; zhou2022optimal; si2024efficient) if predictors are further arranged in a reverse hierarchical order. See Supplementary Material for the detailed illustration.
@@ 2.3 General hierarchical structure
This subsection discusses the case with partially hierarchical predictors and responses, and some of their modes may not be strictly nested within the others, possibly with unknown or even multiple hierarchical relationships among the modes.
We first consider the feature extraction on predictors. A natural way is first to assume a pseudo-full hierarchical structure to predictors, and the proposed method in the previous subsection can then be applied. The pseudo-structure should be consistent with the hierarchical structures intrinsic to the data, and the order of its modes is called an action order. For example, we consider the case with $M=3$, and mode 1 is nested within mode 2. The pseudo-structure can then be with an action order of $\{1,2,3\}$ or $\{3,1,2\}$.
@@
The above proposition proves a nice equivalence of all elements inside $\mathcal{P}_{x}(M)$, however, it is a theoretical result only. In real applications, each action order may have a totally different number of parameters, and a wrongly selected order may end up with a huge number of parameters. As a result, we may choose the one with the smallest size,
@@
The above matrix is no longer orthogonal in general. The philosophy of the above feature extraction is analogous to that of CP decomposition, which concatenates multiple separately extracted features. Note that, comparing with Tucker decomposition, CP decomposition requests much less parameters by sacrificing orthogonality of loading matrices.
@@ 3.1 Ordinary least squares approach
For autoregressive models at (2.12) and (2.13), the coefficient tensor can be denoted by $\mbox{$\mathscr{A}$}_{\mathrm{AR}}\in\mathbb{R}^{q_{1}\times\cdots\times q_{N}\times q_{1}\times\cdots\times q_{N}\times L}$ with $\mbox{$\mathscr{A}$}_{\mathrm{AR}}]_{N}=([\mbox{$\mathscr{A}$}_{1}]_{N},[\mbox{$\mathscr{A}$}_{2}]_{N},\ldots,[\mbox{$\mathscr{A}$}_{L}]_{N})$ being the mode-$N$ sequential matricization. Let $\boldsymbol{\Theta}_{\mathrm{AR}}=[\boldsymbol{\Theta}_{1},\ldots,\boldsymbol{\Theta}_{L}]\in\mathbb{R}^{s\times rL}$, and it holds that $\mbox{$\mathscr{A}$}_{\mathrm{AR}}]_{N}=\boldsymbol{\Lambda}_{y}(\boldsymbol{\Lambda}_{y}^{\top}\boldsymbol{\Lambda}_{y})^{-1}\boldsymbol{\Theta}_{\mathrm{AR}}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})$. Moreover, by reparameterizing $(\boldsymbol{\Lambda}_{y}^{\top}\boldsymbol{\Lambda}_{y})^{-1}\boldsymbol{\Theta}_{\mathrm{AR}}$ into $\boldsymbol{\Theta}_{\mathrm{AR}}$, we have $\mbox{$\mathscr{A}$}_{\mathrm{AR}}]_{N}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}_{\mathrm{AR}}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})$, and then the parameter space can be defined as
@@
$\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{AR}}=\operatorname*{arg\,min}_{\mbox{$\mathscr{A}$}\in\mathcal{S}_{\mathrm{AR}}}\mathcal{L}_{T}(\mbox{$\mathscr{A}$})\quad\text{with}\quad\mathcal{L}_{T}(\mbox{$\mathscr{A}$})=\frac{1}{T}\sum_{t=1}^{T}\left\lVert\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})-[\mbox{$\mathscr{A}$}]_{N}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})\right\rVert_{2}^{2}.$
@@
Denote by $\mbox{$\mathscr{A}$}^{*}$ the true coefficient tensor of the tensor regression or autoregressive model in the previous section, and it depends on the best set of action orders, which is unknown in real applications. In fact, $\mathcal{P}_{x}(M,K_{x})$ and $K_{x}$ in this section are only the running set and numbers of action orders for predictors, respectively, and the situation is the same for $\mathcal{P}_{y}(N,K_{y})$ and $K_{y}$. Fortunately, from Proposition 1, we can project the true coefficient tensor $\mbox{$\mathscr{A}$}^{*}$ into the running sets $\mathcal{P}_{x}(M,K_{x})$ and $\mathcal{P}_{y}(N,K_{y})$, with running numbers $K_{x}$ and $K_{y}$. Such projection has no error, while the number of parameters will be enlarged. Without loss of generality, we will assume $\mathcal{P}_{x}(M,K_{x})$ and $\mathcal{P}_{y}(N,K_{y})$ to be the best sets in establishing theoretical properties, and Section 4 will discuss the selection of these hyperparameters.
3.2 Non-asymptotic properties for tensor regression

### v2

@@ 2.1 Tucker tensor regression and supervised factor modeling
For simplicity, we denote $q_{N+m}=p_{m}$, $s_{N+m}=r_{m}$ and $\boldsymbol{U}_{N+m}=\boldsymbol{V}_{m}$ for $1\leq m\leq M$. Note that the Tucker decomposition at (2.2) is not unique, since $\mbox{$\mathscr{A}$}=\mbox{$\mathscr{G}$}\times_{1}\boldsymbol{U}_{1}\times_{2}\cdots\times_{N+M}\boldsymbol{U}_{N+M}=\left(\mbox{$\mathscr{G}$}\times_{1}\boldsymbol{O}_{1}\times_{2}\cdots\times_{N+M}\boldsymbol{O}_{N+M}\right)\times_{1}\left(\boldsymbol{U}_{1}\boldsymbol{O}_{1}^{-1}\right)\times_{2}\cdots\times_{N+M}\left(\boldsymbol{U}_{N+M}\boldsymbol{O}_{N+M}^{-1}\right)$ for any invertible matrices $\boldsymbol{O}_{i}\in\mathbb{R}^{s_{i}\times s_{i}}$ with $1\leq i\leq N+M$. Without loss of generality, we restrict these factor matrices to be orthonormal, i.e., $\boldsymbol{U}_{j}\in\mathbb{O}^{q_{j}\times s_{j}}$ for each $1\leq j\leq N+M$; see, for example, the higher-order singular value decomposition (HOSVD) in Kolda and Bader (2009). Consequently, the Tucker tensor regression model can be rewritten into
@@
indicating that $\text{col}(\boldsymbol{V})$ contains all useful information of predictors in terms of predicting responses. As a result, define $\boldsymbol{f}_{t}^{\mathrm{response}}=\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t}\times_{n=1}^{N}\boldsymbol{U}_{n}^{\top})=\boldsymbol{U}^{\top}\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})\in\mathbb{R}^{s}$ and $\boldsymbol{f}_{t}^{\mathrm{predictor}}=\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t}\times_{m=1}^{M}\boldsymbol{V}_{m}^{\top})=\boldsymbol{V}^{\top}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})\in\mathbb{R}^{r}$ to be the response and predictor factors, respectively, and then the Tucker tensor regression can be well interpreted from a supervised factor modeling perspective (Wang et al., 2022, 2024; Huang et al., 2025).
The above supervised factor modeling interpretation also motivates us to introduce a new framework to construct a tensor regression model, especially for data with complex structures beyond the capability of Tucker decomposition, or for non-orthogonal loadings in the feature extraction process such as the CP decomposition (Harshman and others, 1970). Specifically, we first extract features or factors from responses and predictors by
@@
respectively. Here, $\mathcal{M}(\cdot):\mathbb{R}^{p_{1}\times\cdots\times p_{M}}\to\mathbb{R}^{r}$ and $\mathcal{N}(\cdot):\mathbb{R}^{q_{1}\times\cdots\times q_{N}}\to\mathbb{R}^{s}$ are feature-extracting operators, with $r$ and $s$ being the number of extracted predictor and response features, respectively, and they do not necessarily belong to the same type. A specific structure is then assumed to the coefficient tensor $\mathscr{A}$ of model (2.1) such that $\boldsymbol{f}_{t}^{\mathrm{response}}=\boldsymbol{\Theta}\boldsymbol{f}_{t}^{\mathrm{predictor}}+\widetilde{\boldsymbol{e}}_{t}$, where $\boldsymbol{\Theta}\in\mathbb{R}^{s\times r}$ and $\widetilde{\boldsymbol{e}}_{t}\in\mathbb{R}^{s}$ are the corresponding coefficient matrix and compressed error term, respectively. To this end, if operators $\mathcal{M}$ and $\mathcal{N}$ are linear, i.e., $\mathcal{M}(\mbox{$\mathscr{X}$}_{t})=\boldsymbol{\Lambda}_{x}^{\top}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})$ and $\mathcal{N}(\mbox{$\mathscr{Y}$}_{t})=\boldsymbol{\Lambda}_{y}^{\top}\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})$, with non-degenerating loading matrices $\boldsymbol{\Lambda}_{x}\in\mathbb{R}^{P\times r}$ and $\boldsymbol{\Lambda}_{y}\in\mathbb{R}^{Q\times s}$, we then can set the coefficient tensor to
@@ 2.2 Purely hierarchical structure
According to technological advances, the available multi-dimensional or tensor-valued data have become more and more complex, and it is very common to observe hierarchical relationships among some modes with their variables being arranged in a similar way. In order to exploit their structural information, we attempt to design a regression model using the proposed framework at (2.4) and (2.5). This subsection starts from the simple case with purely hierarchical responses and predictors.
@@
We first consider the feature extraction on predictors, and assume for simplicity that $M=2$, i.e., $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}}$, and the $p_{1}$ variables of the first mode are nested within each of the $p_{2}$ variables in its second mode. There are two commonly used methods in the literature of factor analysis to extract features from such hierarchical data $\mbox{$\mathscr{X}$}_{t}$: the hierarchical factor model (HFM) in Schmid and Leiman (1957) and higher-order factor model (HOFM) in Yung et al. (1999). The HFM directly extracts features from each level of the hierarchy separately and jointly. In contrast, the HOFM performs feature extraction in a sequential manner along the hierarchy, where features extracted from a lower level are used to summarize those from the upper level. The loading matrices for these two methods are illustrated in Figure 1. Note that the HFM has more than $p_{2}$ features, while the HOFM has $O(rP)$ parameters with $P=p_{1}p_{2}$, making both of them infeasible in high-dimensional settings. Moreover, Tucker decomposition method extracts features simultaneously from all modes, i.e. $\mathcal{M}(\mbox{$\mathscr{X}$}_{t})=\operatorname*{vec}(\boldsymbol{V}_{1}^{\top}\mbox{$\mathscr{X}$}_{t}\boldsymbol{V}_{2})$, and it is not suitable to handle such hierarchical relationships.
@@
Note that the HOFM employs different loadings for each level of the hierarchy. Such flexibility can maintain its performance, while the corresponding number of parameters is also enlarged dramatically for a larger dimension of $M$. However, many recent hierarchical data may have a structure similar to the factorial data, and the variables at each level are also arranged in a similar way. This motivates us to alternatively proposes a sequential feature extraction method with shared loadings within each level of the hierarchy. Specifically, we define the operator $\mathcal{M}(\mbox{$\mathscr{X}$}_{t};\{\boldsymbol{G}_{m}\}_{m=1}^{2})=\boldsymbol{G}_{2}^{\top}(\boldsymbol{I}_{p_{2}}\otimes\boldsymbol{G}_{1}^{\top})\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})$, where $\boldsymbol{G}_{1}\in\mathbb{O}^{p_{1}\times r_{1}}$ and $\boldsymbol{G}_{2}\in\mathbb{O}^{p_{2}r_{1}\times r_{2}}$ are the component matrices for summarizing predictors, and $r_{1}$ and $r_{2}$ are the interim dimensions of features extracted from the corresponding mode. To interpret this construction, we liken the hierarchy of $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}}$ to a tree structure with $p_{2}$ branches, each branch containing $p_{1}$ leaves. For each branch, $\boldsymbol{G}_{1}^{\top}$ extracts $r_{1}$ features from its leaves to represent this branch. Overall, we have $r_{1}p_{2}$ features extracted from all leaves, given by $(\boldsymbol{I}_{p_{2}}\otimes\boldsymbol{G}_{1}^{\top})\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})$. Then the subsequent $\boldsymbol{G}_{2}^{\top}\in\mathbb{R}^{r_{2}\times p_{2}r_{1}}$ combines the extracted features from all branches to form $r_{2}$ features for the entire tree; see Figure 1(c) for the illustration of this hierarchical feature extraction procedure. Therefore, the overall loading matrix connecting the $p_{1}p_{2}$ leaves to the final $r_{2}$ features for the tree is $\boldsymbol{G}_{2}^{\top}(\boldsymbol{I}_{p_{2}}\otimes\boldsymbol{G}_{1}^{\top})$, and the column space of this loading can dwell in an arbitrary subspace of $\mathbb{R}^{r_{2}}$.
@@
Consider a special case with $p_{2}=2$, i.e., predictors $\mbox{$\mathscr{X}$}_{t}$ can be divided into two branches, $\boldsymbol{x}_{1:p_{1},1}$ and $\boldsymbol{x}_{1:p_{1},2}$. Suppose that the first branch has common features $\boldsymbol{G}^{*\top}\boldsymbol{x}_{1:p_{1},1}\in\mathbb{R}^{2}$ and branch-specific one $\boldsymbol{G}_{1}^{*\top}\boldsymbol{x}_{1:p_{1},1}\in\mathbb{R}$, while the second one has the common and branch-specific features of $\boldsymbol{G}^{*\top}\boldsymbol{x}_{1:p_{1},2}\in\mathbb{R}^{2}$ and $\boldsymbol{G}_{2}^{*\top}\boldsymbol{x}_{1:p_{1},2}\in\mathbb{R}^{2}$, respectively. Note that the HOFM uses loading matrices $\boldsymbol{G}_{1}^{(1)}\in\mathbb{R}^{p_{1}\times 3}$ and $\boldsymbol{G}_{1}^{(2)}\in\mathbb{R}^{p_{1}\times 4}$ to extract features from the first and second branches, respectively, while the proposed method employs a common loading matrix $\boldsymbol{G}_{1}\in\mathbb{R}^{p_{1}\times 5}$ to extract features from each branch. Shortly speaking, the HOFM concentrates more on branch-specific features, while our method focuses more on common features. When the data have an irregular structure or one artificially switches the orders of variables at one level, our method may lead to a misleading interpretation, and we may then have to consider the HOFM or even HFM for the feature extracting. As a result, a huge number of parameters will be involved, making it impossible for high-dimensional data.
@@
Consider a special case with $r=s$ and $\boldsymbol{\Theta}$ being diagonal, and then the above coefficient tensor $\mathscr{A}$ will admit a form of tensor train decomposition (Oseledets, 2011; Zhou et al., 2022; Si et al., 2024) if predictors are further arranged in a reverse hierarchical order. See Section B.6 of the Supplementary Material for the detailed illustration.
@@ 2.3 General hierarchical structure
This subsection discusses a more practical case with partially hierarchical predictors and responses, and some of their modes may not be strictly nested within the others, possibly with unknown or even multiple hierarchical relationships among the modes.
We first consider the feature extraction on predictors. A natural way is first to assume a pseudo-full hierarchical structure to predictors, and the proposed method in the previous subsection can then be applied. The pseudo-structure should be consistent with the hierarchical structures intrinsic to the data, and the order of its modes is called an action order. For example, we consider the case with $M=3$, and mode 1 is nested within mode 2. The pseudo-structure can then be represented by action order $(1,2,3)$ or $(3,1,2)$.
@@
The above proposition proves a nice equivalence of all elements inside $\mathcal{P}_{x}(M)$; however, it is a theoretical result only. In real applications, each action order may have a totally different number of parameters, and a wrongly selected order may end up with a huge number of parameters. As a result, we may choose the one with the smallest size,
@@
The above matrix is no longer orthogonal in general. The philosophy of the above feature extraction is analogous to that of CP decomposition, which concatenates multiple separately extracted features. Note that, comparing with Tucker decomposition, CP decomposition requires much less parameters by sacrificing orthogonality of loading matrices.
@@ 3.1 Ordinary least squares approach
For autoregressive models at (2.12) and (2.13), the coefficient tensor can be denoted by $\mbox{$\mathscr{A}$}_{\mathrm{AR}}\in\mathbb{R}^{q_{1}\times\cdots\times q_{N}\times q_{1}\times\cdots\times q_{N}\times L}$ with $\mbox{$\mathscr{A}$}_{\mathrm{AR}}]_{N}=([\mbox{$\mathscr{A}$}_{1}]_{N},[\mbox{$\mathscr{A}$}_{2}]_{N},\ldots,[\mbox{$\mathscr{A}$}_{L}]_{N})$ being the mode-$N$ sequential matricization. Let $\boldsymbol{\Theta}_{\mathrm{AR}}=(\boldsymbol{\Theta}_{1},\ldots,\boldsymbol{\Theta}_{L})\in\mathbb{R}^{s\times rL}$, and it holds that $\mbox{$\mathscr{A}$}_{\mathrm{AR}}]_{N}=\boldsymbol{\Lambda}_{y}(\boldsymbol{\Lambda}_{y}^{\top}\boldsymbol{\Lambda}_{y})^{-1}\boldsymbol{\Theta}_{\mathrm{AR}}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})$. Moreover, by reparameterizing $(\boldsymbol{\Lambda}_{y}^{\top}\boldsymbol{\Lambda}_{y})^{-1}\boldsymbol{\Theta}_{\mathrm{AR}}$ into $\boldsymbol{\Theta}_{\mathrm{AR}}$, we have $\mbox{$\mathscr{A}$}_{\mathrm{AR}}]_{N}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}_{\mathrm{AR}}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})$, and then the parameter space can be defined as
@@
$\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{AR}}=\operatorname*{arg\,min}_{\mbox{$\mathscr{A}$}\in\mathcal{S}_{\mathrm{AR}}}\mathcal{L}_{T}(\mbox{$\mathscr{A}$})\quad\text{with}\quad\mathcal{L}_{T}(\mbox{$\mathscr{A}$})=\frac{1}{T}\sum_{t=1}^{T}\left\lVert\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})-[\mbox{$\mathscr{A}$}]_{N}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})\right\rVert_{2}^{2}.$ (3.1)
@@
Denote by $\mbox{$\mathscr{A}$}^{*}$ the true coefficient tensor of the tensor regression or autoregressive model in the previous section, and it depends on the best set of action orders, which is unknown in real applications. In fact, $\mathcal{P}_{x}(M,K_{x})$ and $K_{x}$ in this section are only the running set and numbers of action orders for predictors, respectively, and the situation is the same for $\mathcal{P}_{y}(N,K_{y})$ and $K_{y}$. Fortunately, from Proposition 1, we can project the true coefficient tensor $\mbox{$\mathscr{A}$}^{*}$ into the running sets $\mathcal{P}_{x}(M,K_{x})$ and $\mathcal{P}_{y}(N,K_{y})$, with running numbers $K_{x}$ and $K_{y}$. Such projection has no error, while the number of parameters will be enlarged. Without loss of generality, we will assume $\mathcal{P}_{x}(M,K_{x})$ and $\mathcal{P}_{y}(N,K_{y})$ to be known in establishing theoretical properties, and Section 4 will discuss the selection of these hyperparameters.
3.2 Non-asymptotic properties for tensor regression settings